# ZipGAN: Super-Resolution-based Generative Adversarial Network Framework for Data Compression of Direct Numerical Simulations

### v2

@@ Abstract
The advancement of high-performance computing has enabled the generation of large direct numerical simulation (DNS) datasets of turbulent flows, driving the need for efficient compression/decompression techniques that reduce storage demands while maintaining fidelity. Traditional methods, such as the discrete wavelet transform, cannot achieve compression ratios of 8 or higher for complex turbulent flows without introducing significant encoding/decoding errors. On the other hand, a super-resolution-based generative adversarial network (SR-GAN), called ZipGAN, can accurately reconstruct fine-scale features, preserving velocity gradients and structural details, even at a compression ratio of 512, thanks to the more efficient representation of the data in compact latent space. Additional benefits are ascribed to adversarial training. The high GAN training time is significantly reduced with a progressive transfer learning approach and, once trained, they can be applied independently of the Reynolds number. It is demonstrated that ZipGAN can enhance dataset temporal resolution without additional simulation overhead by generating high-quality intermediate fields from compressed snapshots. The ZipGAN discriminator can reliably evaluate the quality of decoded fields, ensuring fidelity even in the absence of original DNS fields. Hence, ZipGAN compression/decompression method presents a highly efficient and scalable alternative for large-scale DNS storage and transfer, offering substantial advantages over the DWT methods in terms of compression efficiency, reconstruction fidelity, and temporal resolution enhancement.
@@ 1 Introduction
Lossless compression techniques, such as SZIP [^1] https://docs.hdfgroup.org, are commonly applied to reduce the size of flowfield data, as these methods ensure that the original data can be reconstructed without spurious artifacts. However, the compression ratio achieved by lossless techniques is typically low—usually less than 2—which is often insufficient to compress large-scale flow datasets [3].
Common lossy compression techniques for flowfield data include modal analysis [4], sub-sampling, local re-simulation [5], and wavelet-based methods [6]. These techniques offer significantly higher compression ratios of 4 or more depending on the application. Due to their significantly higher compression, we consider only lossy compression methods in the remainder of this section.
@@ 2 Datasets and data compression methods
Two forced HIT DNS datasets are considered to investigate the effectiveness of the ZipGAN (introduced in Sec. 2.1) framework for the compression of complex turbulent flows. The Taylor-microscale Reynolds numbers of the datasets are $\mathrm{Re_{\lambda}}\approx 90$ (Re90) and $\mathrm{Re_{\lambda}}\approx 350$ (Re350). The domain is discretized on 256 $\times$ 256 $\times$ 256 points for Re90 and 1024 $\times$ 1024 $\times$ 1024 points for Re350. Linear forcing $\textbf{f}=\textit{A}\textbf{u}$ is applied in both simulations, where $A$ is a forcing parameter that is inversely proportional to the eddy turnover time. Simulations are initialized with a synthetic von Karman–Pao spectrum [27]. The minimum cell size, $\mathrm{dx}$, satisfies $\mathrm{\kappa_{\texttt{max}}\eta\geq 1}$, where $\mathrm{\eta}$ is the Kolmogorov length scale and $\mathrm{\kappa_{\texttt{max}}}$ is the maximum discrete wavenumber. The simulation parameters are reported in Table 1. Only the Re90 dataset is used for training the ZipGAN, while the Re350 dataset is reserved for the generalization tests conducted in Sec. 3. Additional details of these datasets can be found in previous works [24, 27].
@@
Re90 $256^{3}$ 90 920 1.98 250 93 $\bullet$ $\bullet$
Re350 $1024^{3}$ 350 16182 1.97 130 3120 $\bullet$
@@
Table 1: Simulation parameters for training and testing datasets. $\mathrm{N}$ is the number of mesh points on each spatial dimension, $\mathrm{Re_{t}}$ is the turbulent Reynolds number, $\mathrm{dx/\eta}$ is the mesh resolution relative to the Kolmogorov length scale $\mathrm{\eta}$, #$\mathrm{\#\,snapshots}$ is the total number of snapshots extracted from each dataset, and $\mathrm{S_{case}}$ is the total dataset size in GB. “Train” and “Test” flags indicate datasets used for model training and testing.
@@
After the Re90 simulation reached a statistically stationary state, 250 snapshots of the three-dimensional velocity vector, u = $(u,v,w)^{T}$ were considered, with five snapshots extracted per eddy-turnover time. To obtain encoded (compressed) fields $\mathrm{\phi_{enc}}$ of different compression ratios, all available snapshots of the instantaneous velocity were filtered using a box filter kernel with discrete downsampling operation. Multiple kernel widths $\mathrm{\Delta=n_{\Delta}\,}\mathrm{dx}$ were used with downsampling factors $\mathrm{n_{\Delta}\in[2,4,8,16]}$. This preprocessing step is denoted an “encoding operation.” The training, validation, and testing datasets are composed of original DNS fields $\mathrm{\phi_{orig}}$ and the corresponding encoded fields $\mathrm{\phi_{enc}}$. The first 190 snapshots of the Re90 dataset were used for training, while the last 60 snapshots were used for validation and testing (training/validation ratio $\approx 0.75$/$0.25$). The validation was then performed on 20 in-sample fields not employed for training. The last 20 snapshots were used to obtain the averaged results discussed in Sec. 3. The time separation between training/validation samples and testing samples was more than 10 eddy-turnover times, thus the two sets of samples are considered statistically uncorrelated.
@@
Loading the entire training dataset into the GPU memory is impractical given memory limitations. To alleviate this, a patch-to-patch training strategy is applied, where sub-domains are randomly extracted from the full domain for each snapshot [28]. The ZipGAN architecture (Sec. 2.1) was trained using $\mathrm{\phi_{enc}}$ and $\mathrm{\phi_{orig}}$ subboxes. Conversely, during the evaluation of testing samples, the entire domain is reconstructed continuously, without employing subboxes as used for training. The velocity components of the original and encoded fields used for training and testing were normalized by the global maximum and minimum of $\mathrm{\phi_{orig}}$ to improve the network’s performance [29]. Training was performed on the Jülich DEEP-EST cluster (DEEP-DAM partition) using four nodes, each containing one NVIDIA V100 32GB GPU, using the TensorFlow framework [30]. The Adam optimizer [31] was utilized with initial learning rate $\alpha_{0}=10^{-4}$ for consistency with previous applications [32, 33].
@@
2.1 Super resolution generative adversarial network architecture (ZipGAN) and training loss function for data compression
@@
This architecture employs a conventional GAN setup for training, consisting of a generator and a discriminator. The generator, employed as a decoder, super-resolves the encoded field $\mathrm{\phi_{enc}}$ to generate the decoded field $\mathrm{\phi_{dec}}$. The discriminator distinguishes between $\mathrm{\phi_{dec}}$ and $\mathrm{\phi_{orig}}$ [34]. As a result, both networks improve their predictions during the adversarial training process. The generator, a CNN based on the SRResNet architecture [35], is designed to recover small-scale features using skip connections and three-dimensional convolutional layers. A central feature of the generator is the residual-in-residual dense block (RRDB) architecture, which includes residual connections and densely connected layer blocks. The generator also includes upsampling layers, which use nearest-neighbor interpolation followed by a convolution layer to enhance the operation. Each upsampling layer doubles the spatial resolution, hence the number of upsampling layers is determined by $\Delta$. The total number of trainable parameters of the generator is approximately 10 million, is mainly driven by the number of filter kernels per convolutional layer (a constant), and is nearly independent of the number of upsampling layers. The discriminator (Fig. 1) is structured as a deep deconvolutional neural network with fully connected and convolutional layers, similar to the generator. It outputs a binary classification and contains around 12 million trainable parameters.
@@
$L_{\mathrm{{adversarial}}}$ $=-\mathbb{E}_{\phi_{\mathrm{{enc}}}\sim p_{\mathrm{enc}}}[\log(\sigma(D(G(\phi_{\mathrm{{enc}}}))-\mathbb{E}_{\phi_{\mathrm{{orig}}}\sim p_{\mathrm{orig}}}[D(\phi_{\mathrm{{orig}}})]))]$
@@
$\mathcal{L}_{\textsc{DISC}}=$ $\mathbb{E}_{\phi_{\mathrm{{orig}}}\sim p_{\mathrm{orig}}}[\log(\sigma(D(\phi_{\mathrm{{orig}}})-\mathbb{E}_{\phi_{\mathrm{{enc}}}\sim p_{\mathrm{enc}}}[D(G(\phi_{\mathrm{{enc}}}))]))]$
@@ 2.2 Discrete wavelet transform method for data compression
For this analysis, the Biorthogonal 4.4 (Bior4.4) wavelet is chosen due to its advantageous properties including symmetry, compact support, and high reconstruction accuracy [13]. The significant coefficients are retained for the reconstruction of the original three-dimensional flow field through the inverse DWT. A detailed explanation of this approach can be found in [13].
@@ 3 Results
Compression method $\mathrm{\theta}$ $\mathrm{n_{\Delta}}$ $\mathrm{S_{\phi_{enc}}}$ $[\mathrm{MB}]$ $\mathrm{N_{\phi_{enc}}}$ ZipGAN $[\mathrm{MB}]$
Re90 Re350 Re90 Re350
@@
DWT ZipGAN 8 $2$ 48 3072 $128^{3}$ $512^{3}$ 42
DWT ZipGAN 64 $4$ 6 410 $64^{3}$ $256^{3}$ 42
DWT ZipGAN 512 $8$ 0.8 51 $32^{3}$ $128^{3}$ 43
DWT ZipGAN 4096 $16$ 0.09 6 $16^{3}$ $64^{3}$ 44
Table 2: Comparison of the encoding/decoding performance using both the DWT and ZipGAN compression methods applied to the Re90 and Re350 datasets. The compression ratio, $\mathrm{\theta}$, is defined as $\mathrm{S_{\phi_{orig}}}/\mathrm{S_{\phi_{enc}}}$, where $S$ represents a field’s size in MB (double precision with three velocity components). $\mathrm{n_{\Delta}}$ is the downsampling factor applied to the original DNS fields $\mathrm{\phi_{orig}}$, $\mathrm{N_{\phi_{enc}}}$ is the number of points in each spatial dimension of the compressed field $\mathrm{\phi_{enc}}$. ”ZipGAN“ refers to the size of the ZipGAN’s generator in MB. Note that for $\mathrm{\theta=2}$ and $\mathrm{\theta=4}$, only the DWT compression method is employed due to the limitations of SR-GAN three-dimensional upsampling layers.
The downsampling factor, $\mathrm{n_{\Delta}}$, of the ZipGAN encoding step increases with $\theta$, yet the size of the ZipGAN remains nearly constant. This is because the overall architecture is preserved, with only the upsampling layers being adjusted. Since this modification has a minimal impact on the number of trainable parameters, the changes to model size are negligible. The actual size of each compressed field, $\mathrm{S_{\phi_{enc}}}$ in MB, is reported for each compression ratio, and is consistent regardless of the compression method used. For simplicity, the corresponding number of points in each spatial dimension of the compressed field, $\mathrm{N_{\phi_{enc}}}$, is provided only for the ZipGAN compression method. However, the total number of points aligns with those in $\mathrm{\phi_{enc}}$ when using the DWT method.
@@ 3.1 Encoding/decoding capabilities for various compression ratios
As velocity-gradient statistics are closely associated with fine-scale motion, data compression/decompression methods are assessed based on their ability to reconstruct these gradients. Figure 2 shows the probability density functions (PDFs) of the normalized velocity gradients for $\mathrm{\phi_{enc}}$, $\mathrm{\phi_{dec}}$, and $\mathrm{\phi_{orig}}$, using the DWT (left) and ZipGAN (right) compression methods on the Re90 dataset. For the DWT, significant deviations from the DNS field are observed for $\mathrm{\theta}>4$, indicating that the quality of the decoded field decreases with $\theta$. In particular, the large distortions observed with the DWT method compared to the original DNS are likely caused by the thresholding process used in wavelet-based compression, which is not based on physical considerations but necessary to achieve the target compression ratio. This leads to the loss of fine-scale turbulence structures and high-frequency components that are critical for DNS representation. As a result, visible distortions are more prominent, especially in regions with sharp gradients or complex structures. Additionally, due to the local wavelet decomposition employed by the DWT, smaller gradients tend to be predicted more accurately than larger ones, which is generally not observed for ZipGAN.
Figure 2: PDF of the normalized velocity gradient of the decoded field, $\mathrm{\phi_{dec}}$, obtained for different compression ratios, $\mathrm{\theta}$, using DWT (left) and ZipGAN (right) compression methods on the Re90 dataset. The input encoded field of the ZipGAN method, $\mathrm{\phi_{enc}}$, is shown for different downsampling factors $\mathrm{n_{\Delta}}$.
@@
Figure 3: 2D slices of instantaneous normalized velocity magnitude and the absolute error of $\phi_{\mathrm{dec,i}}$ fields versus original DNS. The normalized absolute error is given by $\mathrm{\hat{E}=E/\max(E_{\phi_{orig},\phi_{dec,i}},E_{\phi_{orig},\phi_{dec,i}})}$, where $\mathrm{E=\left|\textbf{u}_{\phi_{orig}}-{\textbf{u}}_{\phi_{dec,i}}\right|}$ and the subscript $\mathrm{i}$ indicates either the DWT or the ZipGAN data compression methods.
The decoding capabilities of each compression method are further analyzed in Fig. 4 (left), where additional metrics including the averaged normalized mean squared error (NMSE) of the velocity gradient ¯$\bar{\mathcal{E}}({|\nabla\mathbf{u}|})$ (histogram – lower is better) and the structural similarity index metric (SSIM – higher is better) between $\mathrm{\phi_{dec}}$ and $\mathrm{\phi_{orig}}$ (solid line) are shown. The NMSE quantifies how well the velocity gradients are preserved after decoding, while the SSIM measures the similarity between $\mathrm{\phi_{dec}}$ and $\mathrm{\phi_{orig}}$ fields. The accuracy thresholds are set based on both the averaged NMSE and SSIM for $\mathrm{\theta=8}$, and indicate at what $\mathrm{\theta}$ the decoded field starts to deviate from the original, as shown in Fig. 2 (right). The decoding capabilities are also assessed against a fully supervised SRCNN-based model, called ZipCNN, introduced in Sec. 2.1.
Figure 4: Comparison of the decoding performance at various compression ratios, $\mathrm{\theta}$, based on the averaged NMSE of the velocity gradient, ¯$\bar{\mathcal{E}}({|\nabla\mathbf{u}|})$, (bars), and the SSIM between $\mathrm{\phi_{dec}}$ and $\mathrm{\phi_{orig}}$ (solid line). The comparison is made using DWT, ZipCNN, and ZipGAN compression methods on the Re90 dataset (left) and the Re350 dataset (right).
@@
It is important to note that these findings are independent of the Reynolds number of the HIT dataset. In Fig. 4 (right), the same metrics are analyzed when the ZipGAN and ZipCNN architectures, trained solely on the Re90 dataset, are applied to the Re350 dataset. A consistent $\mathrm{dx/\eta}$ ratio is maintained in this analysis to ensure accurate generalization [29]. The consistency of the findings suggests that retraining on each new dataset may not be required, provided that $\mathrm{dx/\eta}$ is preserved.
For practical, large-scale applications, ZipGAN provides a significant advantage in data compression. The full-scale Re90 dataset is 95,232 MB in size and can be encoded using ZipGAN to 190 MB ($\mathrm{\theta=512}$) plus 43 MB for the encoder’s state. For similar decoding accuracy, the DWT-encoded dataset requires approximately 19,000 MB, which is approximately 100×100$\times$100 × more than the ZipGAN-encoded dataset. A similar compression advantage holds for the larger Re350 dataset, which requires 3,120 GB uncompressed, 6 GB ($\mathrm{\theta=512}$) for ZipGAN, and 780 GB for similar decoding accuracy using the DWT.

### v3

@@ Abstract
The advancement of high-performance computing has enabled the generation of large direct numerical simulation (DNS) datasets of turbulent flows, driving the need for efficient compression/decompression techniques that reduce storage demands while maintaining fidelity. Traditional methods, such as the discrete wavelet transform (DWT), cannot achieve compression ratios of 8 or higher for complex turbulent flows without introducing significant encoding/decoding errors. On the other hand, a super-resolution-based generative adversarial network (SR-GAN), called ZipGAN, can accurately reconstruct fine-scale features, preserving velocity gradients and structural details, even at a compression ratio of 512, thanks to the more efficient representation of the data in a compact latent space. Additional benefits are ascribed to adversarial training. The high GAN training time is significantly reduced with a progressive transfer learning approach and, once trained, they can be applied independently of the Reynolds number. It is demonstrated that ZipGAN can enhance dataset temporal resolution without additional simulation overhead by generating high-quality intermediate fields from compressed snapshots. The ZipGAN discriminator can reliably evaluate the quality of decoded fields, ensuring fidelity even in the absence of original DNS fields. Hence, ZipGAN compression/decompression method presents a highly efficient and scalable alternative for large-scale DNS storage and transfer, offering substantial advantages over the DWT methods in terms of compression efficiency, reconstruction fidelity, and temporal resolution enhancement.
@@ 1 Introduction
Lossless compression techniques, such as SZIP [^1] https://support.hdfgroup.org, are commonly applied to reduce the size of flow field data, as these methods ensure that the original data can be reconstructed without spurious artifacts. However, the compression ratio achieved by lossless techniques is typically low—usually less than 2—which is often insufficient to compress large-scale flow datasets [3].
Common lossy compression techniques for flow field data include modal analysis [4], sub-sampling, local re-simulation [5], and wavelet-based methods [6]. These techniques offer significantly higher compression ratios of 4 or more, depending on the application. Due to their significantly higher compression, we consider only lossy compression methods in the remainder of this section.
@@ 2 Datasets and data compression methods
Two forced HIT DNS datasets are considered to investigate the effectiveness of the ZipGAN framework (introduced in Sec. 2.1) for the compression of complex turbulent flows. The Taylor-microscale Reynolds numbers of the datasets are $\mathrm{Re_{\lambda}}\approx 90$ (Re90) and $\mathrm{Re_{\lambda}}\approx 210$ (Re210). The domain is discretized on 256 $\times$ 256 $\times$ 256 points for Re90 and 1024 $\times$ 1024 $\times$ 1024 points for Re210. Linear forcing $\textbf{f}=\textit{A}\textbf{u}$ is applied in both simulations, where $A$ is a forcing parameter that is inversely proportional to the eddy turnover time. Simulations are initialized with a synthetic von Kármán–Pao spectrum [27]. The minimum cell size, $\mathrm{dx}$, satisfies $\mathrm{\kappa_{\texttt{max}}\eta\geq 1}$, where $\mathrm{\eta}$ is the Kolmogorov length scale and $\mathrm{\kappa_{\texttt{max}}}$ is the maximum discrete wavenumber. The simulation parameters are reported in Table 1. Only the Re90 dataset is used for training the ZipGAN, while the Re210 dataset is reserved for the generalization tests conducted in Sec. 3. Additional details of these datasets can be found in previous works [24, 27].
@@
Re90 $256^{3}$ 90 455 1.98 250 93 $\bullet$ $\bullet$
Re210 $1024^{3}$ 210 1300 1.97 130 3120 $\bullet$
@@
Table 1: Simulation parameters for training and testing datasets. $\mathrm{N}$ is the number of mesh points on each spatial dimension, $\mathrm{Re_{t}}$ is the turbulent Reynolds number, $\mathrm{dx/\eta}$ is the mesh resolution relative to the Kolmogorov length scale $\mathrm{\eta}$, $\mathrm{\#\,snapshots}$ is the total number of snapshots extracted from each dataset, and $\mathrm{S_{case}}$ is the total dataset size in GB. “Train” and “Test” flags indicate datasets used for model training and testing.
@@
After the Re90 simulation reached a statistically stationary state, 250 snapshots of the three-dimensional velocity vector, u = $(u,v,w)^{T}$ were considered, with five snapshots extracted per eddy-turnover time. To obtain encoded (compressed) fields $\mathrm{\phi_{enc}}$ of different compression ratios, all available snapshots of the instantaneous velocity were filtered using a box filter kernel with a discrete downsampling operation. Multiple kernel widths $\mathrm{\Delta=n_{\Delta}\,}\mathrm{dx}$ were used with downsampling factors $\mathrm{n_{\Delta}\in[2,4,8,16]}$. This preprocessing step is denoted as an “encoding operation.” The training, validation, and testing datasets are composed of original DNS fields $\mathrm{\phi_{orig}}$ and the corresponding encoded fields $\mathrm{\phi_{enc}}$. The first 190 snapshots of the Re90 dataset were used for training, while the last 60 snapshots were used for validation and testing (training/validation ratio $\approx 0.75$/$0.25$). The validation was then performed on 20 in-sample fields not employed for training. The last 20 snapshots were used to obtain the averaged results discussed in Sec. 3. The time separation between training/validation samples and testing samples was more than 10 eddy-turnover times, thus, the two sets of samples are considered statistically uncorrelated.
@@
Loading the entire training dataset into GPU memory is impractical given memory limitations. To alleviate this, a patch-to-patch training strategy is applied, where sub-domains are randomly extracted from the full domain for each snapshot [28]. The ZipGAN framework (Sec. 2.1) was trained using $\mathrm{\phi_{enc}}$ and $\mathrm{\phi_{orig}}$ subboxes. Conversely, during the evaluation of testing samples, the entire domain is reconstructed continuously, without employing subboxes as used for training. The velocity components of the original and encoded fields used for training and testing were normalized by the global maximum and minimum of $\mathrm{\phi_{orig}}$ to improve the network’s performance [29]. Training was performed on the Jülich DEEP-EST cluster (DEEP-DAM partition) using four nodes, each containing one NVIDIA V100 32GB GPU, using the TensorFlow framework [30]. The Adam optimizer [31] was utilized with initial learning rate $\alpha_{0}=10^{-4}$ for consistency with previous applications [32, 33].
@@
2.1 Super resolution generative adversarial network architecture (ZipGAN) and training loss function for data decompression
@@
This architecture employs a conventional GAN setup for training, consisting of a generator and a discriminator. The generator, employed as a decoder, super-resolves the encoded field $\mathrm{\phi_{enc}}$ to generate the decoded field $\mathrm{\phi_{dec}}$. The discriminator distinguishes between $\mathrm{\phi_{dec}}$ and $\mathrm{\phi_{orig}}$ [34]. As a result, both networks improve their predictions during the adversarial training process. The generator, a CNN based on the SRResNet architecture [35], is designed to recover small-scale features using skip connections and three-dimensional convolutional layers. A central feature of the generator is the residual-in-residual dense block (RRDB) architecture, which includes residual connections and densely connected layer blocks. The generator also includes upsampling layers, which use nearest-neighbor interpolation followed by a convolution layer to enhance the operation. Each upsampling layer doubles the spatial resolution, hence, the number of upsampling layers is determined by $\Delta$. The total number of trainable parameters of the generator is approximately 10 million, and is mainly driven by the number of filter kernels per convolutional layer (a constant), and is nearly independent of the number of upsampling layers. The discriminator (Fig. 1) is structured as a deep deconvolutional neural network with fully connected and convolutional layers, similar to the generator. It outputs a binary classification and contains around 12 million trainable parameters.
@@
$L_{\mathrm{{adversarial}}}$ $=-\mathbb{E}_{\phi_{\mathrm{{enc}}}\sim p_{\mathrm{enc}}}[\log(\sigma(D(G(\phi_{\mathrm{{enc}}}))-\mathbb{E}_{\phi_{\mathrm{{orig}}}\sim p_{\mathrm{orig}}}[D(\phi_{\mathrm{{orig}}})]))]\$
@@
$\mathcal{L}_{\textsc{DISC}}=$ $\mathbb{E}_{\phi_{\mathrm{{orig}}}\sim p_{\mathrm{orig}}}[\log(\sigma(D(\phi_{\mathrm{{orig}}})-\mathbb{E}_{\phi_{\mathrm{{enc}}}\sim p_{\mathrm{enc}}}[D(G(\phi_{\mathrm{{enc}}}))]))]\$
@@ 2.2 Discrete wavelet transform method for data compression
For this analysis, the Biorthogonal 4.4 (Bior4.4) wavelet is chosen due to its advantageous properties, including symmetry, compact support, and high reconstruction accuracy [13]. The significant coefficients are retained for the reconstruction of the original three-dimensional flow field through the inverse DWT. A detailed explanation of this approach can be found in [13].
@@ 3 Results
Compression
method
$\mathrm{\theta}$ $\mathrm{n_{\Delta}}$
$\mathrm{S_{\phi_{enc}}}$ $\mathrm{MB}]$
$\mathrm{N_{\phi_{enc}}}$
ZipGAN $\mathrm{MB}]$
Re90 Re210 Re90 Re210
@@
DWT
ZipGAN
8
22
48 3072
$128^{3}$
$512^{3}$
42
DWT
ZipGAN
64
44
6 410
$64^{3}$
$256^{3}$
42
DWT
ZipGAN
512
88
0.8 51
$32^{3}$
$128^{3}$
43
DWT
ZipGAN
4096
1616
0.09 6
$16^{3}$
$64^{3}$
44
Table 2: Comparison of the encoding/decoding performance using both the DWT and ZipGAN compression methods applied to the Re90 and Re210 datasets. The compression ratio, $\mathrm{\theta}$, is defined as $\mathrm{S_{\phi_{orig}}}/\mathrm{S_{\phi_{enc}}}$, where $S$ represents a field’s size in MB (double precision with three velocity components). $\mathrm{n_{\Delta}}$ is the downsampling factor applied to the original DNS fields $\mathrm{\phi_{orig}}$, $\mathrm{N_{\phi_{enc}}}$ is the number of points in each spatial dimension of the compressed field $\mathrm{\phi_{enc}}$. ”ZipGAN“ refers to the size of the ZipGAN’s generator in MB. Note that for $\mathrm{\theta=2}$ and $\mathrm{\theta=4}$, only the DWT compression method is employed due to the limitations of SR-GAN three-dimensional upsampling layers.
The downsampling factor, $\mathrm{n_{\Delta}}$, of the ZipGAN encoding step increases with $\theta$, yet the size of the ZipGAN architecture remains nearly constant. This is because the overall architecture is preserved, with only the upsampling layers being adjusted. Since this modification has a minimal impact on the number of trainable parameters, the changes to model size are negligible. The actual size of each compressed field, $\mathrm{S_{\phi_{enc}}}$ in MB, is reported for each compression ratio, and is consistent regardless of the compression method used. For simplicity, the corresponding number of points in each spatial dimension of the compressed field, $\mathrm{N_{\phi_{enc}}}$, is provided only for the ZipGAN compression method. However, the total number of points aligns with those in $\mathrm{\phi_{enc}}$ when using the DWT method.
@@ 3.1 Encoding/decoding capabilities for various compression ratios
As velocity-gradient statistics are closely associated with fine-scale motion, data compression/decompression methods are assessed based on their ability to reconstruct these gradients. Figure 2 shows the probability density functions (PDFs) of the normalized velocity gradients for $\mathrm{\phi_{enc}}$, $\mathrm{\phi_{dec}}$, and $\mathrm{\phi_{orig}}$, using the DWT (left) and ZipGAN (right) compression methods on the Re90 dataset. For the DWT, significant deviations from the DNS field are observed for $\mathrm{\theta}>4$, indicating that the quality of the decoded field decreases with $\theta$. In particular, the large distortions observed with the DWT method compared to the original DNS are likely caused by the thresholding process used in wavelet-based compression, which is not based on physical considerations but is necessary to achieve the target compression ratio. This leads to the loss of fine-scale turbulence structures and high-frequency components that are critical for DNS representation. As a result, visible distortions are more prominent, especially in regions with sharp gradients or complex structures. Additionally, due to the local wavelet decomposition employed by the DWT, smaller gradients tend to be predicted more accurately than larger ones, which is generally not observed for ZipGAN.
Figure 2: PDF of the normalized velocity gradient of the decoded field, $\mathrm{\phi_{dec}}$, obtained for different compression ratios, $\mathrm{\theta}$, using DWT (left) and ZipGAN (right) compression methods on the Re90 dataset. The input encoded field of the ZipGAN method, $\mathrm{\phi_{enc}}$, is shown for different downsampling factors, $\mathrm{n_{\Delta}}$.
Figure 3: 2D slices of instantaneous normalized velocity magnitude and the absolute error of $\phi_{\mathrm{dec,i}}$ fields versus original DNS. The normalized absolute error is given by $\mathrm{\hat{E}=E/\max(E_{\phi_{orig},\phi_{dec,i}},E_{\phi_{orig},\phi_{dec,i}})}$, where $\mathrm{E=\left|\textbf{u}_{\phi_{\mathrm{orig}}}-{\textbf{u}}_{\phi_{\mathrm{dec,i}}}\right|}$ and the subscript $\mathrm{i}$ indicates either the DWT or the ZipGAN data compression methods.
@@
The decoding capabilities of each compression method are further analyzed in Fig. 4 (left), where additional metrics including the averaged normalized mean squared error (NMSE) of the velocity gradient $\bar{\mathcal{E}}({|\nabla\mathbf{u}|})$ (histogram – lower is better) and the structural similarity index metric (SSIM – higher is better) between $\mathrm{\phi_{dec}}$ and $\mathrm{\phi_{orig}}$ (solid line) are shown. The NMSE quantifies how well the velocity gradients are preserved after decoding, while the SSIM measures the similarity between $\mathrm{\phi_{dec}}$ and $\mathrm{\phi_{orig}}$ fields. The accuracy thresholds are set based on both the averaged NMSE and SSIM for $\mathrm{\theta=8}$, and indicate at what $\mathrm{\theta}$ the decoded field starts to deviate from the original, as shown in Fig. 2 (right). The decoding capabilities are also assessed against a fully supervised SRCNN-based model, called ZipCNN, introduced in Sec. 2.1.
Figure 4: Comparison of the decoding performance at various compression ratios, $\mathrm{\theta}$, based on the averaged NMSE of the velocity gradient, $\bar{\mathcal{E}}({|\nabla\mathbf{u}|})$, (bars), and the SSIM between $\mathrm{\phi_{dec}}$ and $\mathrm{\phi_{orig}}$ (solid line). The comparison is made using DWT, ZipCNN, and ZipGAN compression methods on the Re90 dataset (left) and the Re210 dataset (right).
@@
It is important to note that these findings are independent of the Reynolds number of the HIT dataset. In Fig. 4 (right), the same metrics are analyzed when the ZipGAN and ZipCNN architectures, trained solely on the Re90 dataset, are applied to the Re210 dataset. A consistent $\mathrm{dx/\eta}$ ratio is maintained in this analysis to ensure accurate generalization [29]. The consistency of the findings suggests that retraining on each new dataset may not be required, provided that $\mathrm{dx/\eta}$ is preserved.
For practical, large-scale applications, ZipGAN provides a significant advantage in data compression. The full-scale Re90 dataset is 95,232 MB in size and can be encoded using ZipGAN to 190 MB ($\mathrm{\theta=512}$) plus 43 MB for the encoder’s state. For similar decoding accuracy, the DWT-encoded dataset requires approximately 19,000 MB, which is approximately $100\times$ more than the ZipGAN-encoded dataset. A similar compression advantage holds for the larger Re210 dataset, which requires 3,120 GB uncompressed, 6 GB ($\mathrm{\theta=512}$) for ZipGAN, and 780 GB for similar decoding accuracy using the DWT.